# ———————————————————————————————————–

### v2

@@ Abstract
We propose a two step strategy for estimating one-dimensional dynamical parameters of a quantum Markov chain, which involves quantum post-processing the output using a coherent quantum absorber and a “pattern counting” estimator computed as a simple additive functional of the outcomes trajectory produced by sequential, identical measurements on the output units. We provide strong theoretical and numerical evidence that the estimator achieves the quantum Cramé-Rao bound in the limit of large output size.
@@
Our estimation method is underpinned by an asymptotic theory of translationally invariant modes (TIMs) built as averages of shifted tensor products of output operators, labelled by binary patterns. For large times, the TIMs form a bosonic algebra and the output state approaches a joint coherent state of the TIMs whose amplitude depends linearly on the mismatch between system and absorber parameters. Moreover, in the asymptotic regime the TIMs capture the full quantum Fisher information of the output state. While directly probing the TIMs’ quadratures seems impractical, we show that the standard sequential measurement is an effective joint measurement of all the TIMs number operators; indeed, we show that counts of different binary patterns extracted from the measurement trajectory have the expected joint Poisson distribution. Together with the displaced-null methodology of Girotti et al. (2024) this provides a computationally efficient estimator which only depends on the total number of patterns. This opens the way for similar estimation strategies in continuous-time dynamics, expanding the results of Yang et al. (2023).
I Introduction
Quantum statistical inference Helstrom (1976); Holevo (2011); Hayashi (2005); Paris (2008); Tóth and Apellaniz (2014); Demkowicz-Dobrzański et al. (2020); Banaszek et al. (2013); Albarelli et al. (2020); Sidhu and Kok (2020) provides the mathematical framework for enhanced metrology Giovannetti et al. (2005); Fujiwara and Imai (2008); Giovannetti et al. (2011); Escher et al. (2011); Demkowicz-Dobrzański et al. (2012); Girolami et al. (2014); Smirne et al. (2016); Seveso et al. (2017); Haase et al. (2018); Pang and Brun (2014); Rossi et al. (2020); Zhou and Jiang (2021), imaging Tsang et al. (2016); Tsang (2021); Lupo et al. (2020); Fiderer et al. (2021); Oh et al. (2021), waveform and noise estimation Tsang et al. (2011); Berry et al. (2015); Ng et al. (2016); Norris et al. (2016); Shi and Zhuang (2023); Sung et al. (2019); Tsang (2023), and quantum sensing applications Dowling and Seshadreesan (2015); Degen et al. (2017); Pezzè et al. (2018); Marciniak et al. (2022); Zwick and Álvarez (2023) including time keeping Robinson et al. (2024), magnetometry Jones et al. (2009); Amorós-Binefa and Kołodyński (2021); Brask et al. (2015); Albarelli et al. (2017), biomedical sensing Aslam et al. (2023), thermometry Correa et al. (2015); Mehboudi et al. (2019), gravitational wave detection Abbott et al. (2016); Tse et al. (2019); Bailes et al. (2021); Christensen and Meyer (2022).
@@ I Introduction
The cornerstone of quantum estimation is the quantum Cramér-Rao bound (QCRB) Holevo (2011); Helstrom (1976); Belavkin (1976); Braunstein and Caves (1994) which places a fundamental restriction on the precision in estimating unknown parameters of a quantum state. For one-dimensional parameters the bound is attainable in the limit of many copies, by measuring a specific observable called the symmetric logarithmic derivative, which is the quantum analogue of the classical score function. However, when dealing with complex models involving correlated states of many-body systems, the optimal measurement may be hard to compute and implement in practice, and in the case of pure states (and more generally, rank deficient states) such a measurement is highly non-unique. Therefore, it is particularly important to devise realistic measurement schemes which allow the estimation of unknown parameters with close to optimal precision, by means of computationally efficient estimators.
@@
In this paper we provide a general measurement and data processing protocol for estimating an arbitrary dynamical parameter of a quantum Markov chain (QMC), which satisfies the above criteria. A QMC is a discrete time model of an open quantum system, in which the system interacts successively with a sequence of identically prepared “noise units” representing the environment, cf. Figure 1 a). The setup is similar to Haroche’s photon-box one-atom maser Haroche and Raimond (2006) and to that used in quantum collision models Ciccarello et al. (2022), and provides a physical mechanism for generating versatile many-body states such as matrix product states Perez-Garcia et al. (2007); Schön et al. (2005) and finitely correlated states Fannes et al. (1992, 1994). By discretising time, QMCs can be used to model continuous-time dynamics of a Markovian open system coupled with Bosonic input-output channels Attal and Pautrat (2006a); Verstraete and Cirac (2010); Gough (2004).
@@
Figure 1: Basic elements of the pattern counting estimator. Panel a) A quantum Markov chain as a system interacting sequentially with the environment via a parameter dependent unitary $U_{\theta}$. The first stage estimator $\tilde{\theta}_{n}$ is obtained by performing a standard sequential measurement on the output and equating empirical and expected counts. Panel b) Post-processing the output using a coherent absorber. When system and absorber parameters match, the output is identical to the input (vacuum) Panel c) After the first estimation stage the absorber is fixed at a value $\theta_{\rm abs}=\tilde{\theta}_{n}-\delta_{n}$ where $\tilde{\theta}_{n}$ is the preliminary estimator and $\delta_{n}$ is the parameter shift required by the displaced-null measurement theory Girotti et al. (2024). The output generated by the system and absorber dynamics with unitary $V_{\theta_{\rm abs}}U_{\theta}$ is measured sequentially in the standard basis. Panel d) Given a measurement trajectory, excitation patterns are identified as binary sequences starting and ending with a $1$ separated by long sequences of 00s. The final estimator is a correction to the preliminary estimator which depends only on the total number of patterns $\sum_{\alpha}N_{\alpha,n}$, the QFI $f$ at $\tilde{\theta}_{n}$ and the displacement parameter $\tau_{n}$.
@@
For clarity, it is useful to distinguish between two mainstream approaches to parameter estimation in quantum open systems. In the setting of Huelga et al. (1997); Benatti et al. (2014); Smirne et al. (2016); Sekatski et al. (2017); Demkowicz-Dobrzański et al. (2017); Zhou et al. (2018); Górecki et al. (2020), the quantum system undergoes a noisy evolution depending on an unknown parameter, and the experimenter tries to extract information about the parameter by repeatedly applying instantaneous direct measurements and control operations while the system is evolving. In contrast, the setting adopted in this paper is commonly used in quantum optics and input-output theory Gardiner and Zoller (2004); Wiseman and Milburn (2010); Gough and James (2009); Combes et al. (2017) where the experimenter does not have direct access to the system but can measure the output field of an environment channel coupled to the system. This allows the experimenter to track the conditional state of the system by means of stochastic filtering equations (Belavkin, 1994; Dalibard et al., 1992; Wiseman and Milburn, 1993; Bouten et al., 2007; Carmichael, 1993) and control it using feedback. As these techniques require full knowledge of the system’s dynamical parameters, it is important to devise tools for estimating such parameters from the stochastic trajectory of the measurement record. Since the early works Mabuchi (1996); Gambetta and Wiseman (2001), many aspects of continuous-time estimation have been investigated, including adaptive estimation Berry and Wiseman (2002); Pope et al. (2004) filtering methods Ralph et al. (2011); Chase and Geremia (2009); Six et al. (2015), Heisenberg scaling (Macieszczak et al., 2016; Albarelli et al., 2017, 2018; Ilias et al., 2022), sensing with error correction (Plenio and Huelga, 2016), Bayesian estimation (Gammelmark and Mølmer, 2013; Negretti and Mølmer, 2013; Kiilerich and Mølmer, 2016; Ralph et al., 2017; Zhang et al., 2019), quantum smoothing (Tsang, 2009a, b, 2010; Guevara and Wiseman, 2015), estimation of linear systems (Guţǎ and Yamamoto, 2016; Genoni, 2017; Levitt and Guţă, 2017; Levitt et al., 2018), central limit and large deviations theory for trajectories Garrahan and Lesanovsky (2010); van Horssen and Guţă (2015); Burgarth et al. (2015); Carollo et al. (2019), concentration bounds for time averaged observables Benoist et al. (2022); Girotti et al. (2023), estimation with feedback control Fallani et al. (2022).
However, a major problem in this area has been that standard measurement protocols such as counting and homodyne do not achieve the ultimate precision limit prescribed by the QCRB in terms of the quantum Fisher information (QFI) of the output state Guţă (2011); Gammelmark and Mølmer (2014); Catana et al. (2015); Guţă and Kiukas (2015, 2017); Genoni (2017).
@@
Two recent papers Godley and Guţă (2023); Yang et al. (2023) have addressed this problem by introducing the idea of quantum post-processing of the output state using a quantum coherent absorber Stannigel et al. (2012). For a given QMC reference dynamics, the absorber takes the system’s output as its own input and is characterised by the property that it “reverts” the action of the system, so that its output is a trivial product state (vacuum in continuous-time dynamics), cf. Figure 1 b). In the statistical estimation framework, the absorber is set to a particular reference parameter of the QMC dynamics, so that a small deviation of the true parameter from this value will lead to non-trivial output statistics which can be used to estimate the deviation as illustrated in Figure 1 c). In Godley and Guţă (2023), two of the present authors proved that the QCRB can be achieved by performing sequential, adaptive measurements on the output units, after the interaction with the coherent absorber. In addition, the adaptive measurement can be implemented efficiently in a Markovian fashion. The work Yang et al. (2023) deals with the same problem but in the continuous-time setting, and proposes to perform a standard counting measurement (instead of an adaptive one) in conjunction with post-processing using the coherent absorber. In both papers the final estimator was computed from the measurement trajectory using the maximum likelihood method, which becomes inefficient for long trajectories.
@@
The strategy proposed here is similar to that of Yang et al. (2023), albeit in discrete rather than continuous time, but strengthens it in several important aspects. Firstly, we employ the technique of displaced-null measurements Girotti et al. (2024) to provide a precise recipe for choosing the coherent absorber parameter. As we explain below, this is an important technical detail, as the intuitive choice of parameters fails to achieve the QCRB in the limit of long trajectories. Secondly, our strategy employs a two step adaptive procedure which allows us to compute the final estimator as a simple linear transformation of the total number of “pattern counts” which can be easily extracted from the measurement trajectory, cf. Figure 1 d). This circumvents the computational issues associated with the maximum likelihood estimator. Thirdly, we provide strong theoretical evidence that the final estimator achieves the QCRB in the limit of large times. This is based on a novel representation of the output in terms of translationally invariant modes which are shown to satisfy the quantum local asymptotic normality property Guţă and Kahn (2006); Kahn and Guţă (2009); Guţă et al. (2008); Butucea et al. (2016).
@@
We now give a brief summary of the two steps estimation strategy proposed here and the related mathematical results. We consider a QMS whose dynamics depends on a one-dimensional parameter $\theta$ which we aim to estimate by measuring the output state produced after $n$ time steps. In the first stage we run the QMC dynamics with the unknown parameter $\theta$ for $1\ll\tilde{n}\ll n$ time steps and measure the noise units in a fixed basis, cf. Figure 1 a). From the total counts statistics we construct a rough estimator $\tilde{\theta}_{n}$ by matching the empirical frequency to its expected value. In general this estimator is not optimal but its mean square error has the standard $1/\tilde{n}$ scaling (Girotti et al. (2023)). In the second stage we run the system and absorber QMC for the remaining $n-\tilde{n}$ time steps and measure the output in the standard basis. If the absorber parameter matched the true system parameter $\theta$, this measurement would produce a string of 00s (corresponding to no counts in continuous-time) cf. Figure 1 b). Therefore, it would seem natural to choose the absorber parameter to be $\tilde{\theta}_{n}$, our best guess at the unknown parameter $\theta$. However, this choice is unsuitable since for small deviations $\Delta_{n}=\theta-\tilde{\theta}_{n}$, the counting statistics depends quadratically on $\Delta_{n}$, which prevents the estimation of $\theta$ at standard $1/n$ rate. This non-identifiability issue is explained in detail in Girotti et al. (2024), which also provides the solution to this problem. We deliberately set the absorber parameter at $\theta_{\rm abs}=\tilde{\theta}_{n}-\delta_{n}$, which is away from the best guess by a small “displacement” $\delta_{n}\downarrow 0$ chosen to be larger than the uncertainty $|\Delta_{n}|$. This allows us to unambiguously identify $\theta$ from counts statistics. Stage two of the estimation procedure is illustrated in Figure 1 c).
@@
We turn now to the question of estimating $\theta$ from the counts trajectory $\omega=(\omega_{1},\omega_{2},\dots,\omega_{n})$ of the second stage measurement (setting $n-\tilde{n}$ to $n$ for simplicity). Since $\theta-\theta_{\rm abs}$ is vanishingly small (even with the extra displacement), $\omega$ will typically contain a small number of $1$s while most of the outcomes will be 00, cf. Figure 1 d). This allows us split the trajectory into long sequences of 00s and in between them, binary “excitation patterns” starting and ending with a $1$. For each pattern $\alpha$ (e.g. $1$, $11$, $101$ etc.) we count the number of occurrences $N_{\alpha,n}$. In Theorem 3 we show that in the limit of large $n$ the counts $N_{\alpha,n}$ become independent Poisson variables whose intensities are $\lambda_{\alpha}u^{2}$ where $u=\sqrt{n}(\theta-\tilde{\theta}_{\rm abs})$ is the “local parameter” and $\lambda_{\alpha}$ is a model dependent coefficient. Moreover the total Fisher information of the Poisson variables is equal to the output QFI, which shows that the pattern counts statistics capture the full information of the output state. Using this asymptotic behaviour, we construct a simple estimator $\hat{\theta}_{n}$ (cf. equations (23) and (22)) which is linear in the total pattern count, and we argue why it should achieve the QCRB in the limit of large $n$. To summarise, the two step procedure provides a computationally and statistically efficient estimation method which involves only standard basis measurements and a minimal amount of “quantum post-processing” implemented by the coherent absorber.
@@
For a more in-depth understanding of why the excitation pattern counts have asymptotically Poisson distributions, we refer to sections IV and V where we develop a theory of translationally invariant modes (TIMs) of the output. These modes turn out to capture all statistical information about the unknown parameter, and can be measured simultaneously and optimally by performing the sequential standard output measurement. For each excitation pattern $\alpha=(\alpha_{1},\dots,\alpha_{k})\in\{0,1\}^{k}$ we define the creation operator $A^{*}_{\alpha}(n)$ on the output chain of length $n$. This consists of a running average
@@
where $\sigma^{\alpha}_{i}$ is the tensor product of the the type $\sigma^{\alpha}=\sigma^{\alpha_{1}}\otimes\dots\otimes\sigma^{\alpha_{k}}$ where $\sigma^{0}=\mathbf{1}$ and $\sigma^{1}=\sigma^{+}=|1\rangle\langle 0|$, with first tensor acting on position $i$ of the output chain. In Proposition 1 and Corollary 1 we show that asymptotically with $n$, by applying the creation operators to the reference (vacuum) state $|0\rangle^{\otimes n}$, we obtain Fock-type states with different excitation pattern numbers. The creation and annihilation operators satisfy the Bosonic commutation relations with each excitation pattern being an independent mode. For large $n$, a Fock state is a superposition of basis states consisting of long sequences of 00s interspersed with the corresponding patterns appearing in any possible order. One of our key results, Theorem 2 shows that when the gap between system and absorber parameters scales as $\theta-\theta_{\rm abs}=u/\sqrt{n}$, the quadratures of the excitation pattern modes satisfy the Central Limit Theorem and the corresponding joint state is a product of coherent states whose amplitudes are linear in the local parameter $u$. The total QFI of this multimode coherent state is equal to the output QFI, showing that the TIMs contain all statistical information about the dynamics. We also prove separately, that the number operators of the TIMs have asymptotic Poisson distributions, as expected for a coherent state. Together with the result of Theorem 3, this completes a circle of ideas, which played a crucial role in formulating our estimation strategy. In a nutshell, when looking at the output from the perspective of the TIMs, one deals with a simple Gaussian estimation problem. Using the displaced-null method, we can achieve the QCRB by measuring the number operators of the TIMs and such a measurement can be implemented by simple sequential counting measurements followed by the extraction of pattern counts from the measurement trajectory $\omega$.
@@
The paper is organised as follows. In section II we give a brief review of quantum estimation theory and the displaced-null measurement technique developed in Girotti et al. (2024). In section III we introduce the notion of QMC and the estimation problem, together with the idea of quantum post-processing using a coherent absorber. In section IV we define the translationally invariant modes of the output and establish their Fock space properties. In section V we show that the restriction of the output state to the TIMs is a coherent state whose amplitude is linear in the local parameter and whose QFI is equal to the output QFI (cf. Theorem 2 and Corollary 2). In section VI we establish that the excitation pattern counts obtained from the sequential output measurement have asymptotically Poisson distribution (cf. Theorem 3). In section VII we formulate our measurement and estimation strategy and define the ”pattern counts” estimator. Finally in section VIII we present results of a simulation study confirming the earlier theoretical results.
II Quantum estimation and the displaced null measurement technique

### v3

@@ Abstract
We propose a two step strategy for estimating one-dimensional dynamical parameters of a quantum Markov chain, which involves quantum post-processing the output using a coherent quantum absorber and a “pattern counting” estimator computed as a simple additive functional of the outcomes trajectory produced by sequential, identical measurements on the output units. We provide strong theoretical and numerical evidence that the estimator achieves the quantum Cramér-Rao bound in the limit of large output size.
@@
Our estimation method is underpinned by an asymptotic theory of translationally invariant modes (TIMs) built as averages of shifted tensor products of output operators, labelled by binary patterns. For large times, the TIMs form a bosonic algebra and the output state approaches a joint coherent state of the TIMs, whose amplitude depends linearly on the mismatch between system and absorber parameters. Moreover, in the asymptotic regime, the TIMs capture the full quantum Fisher information of the output state. While directly probing the TIMs’ quadratures seems impractical, we show that the standard sequential measurement is an effective joint measurement of all the TIMs number operators; indeed, we show that counts of different binary patterns extracted from the measurement trajectory have the expected joint Poisson distribution. Together with the displaced-null methodology of [1] this provides a computationally efficient estimator which only depends on the total number of patterns. This opens the way for similar estimation strategies in continuous-time dynamics, expanding the results of [2].
1 Introduction
Quantum statistical inference [3, 4, 5, 6, 7, 8, 9, 10, 11] provides the mathematical framework for enhanced metrology [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23], imaging [24, 25, 26, 27, 28], waveform and noise estimation [29, 30, 31, 32, 33, 34, 35], and quantum sensing applications [36, 37, 38, 39, 40] including time keeping [41], magnetometry [42, 43, 44, 45], biomedical sensing [46], thermometry [47, 48], gravitational wave detection [49, 50, 51, 52].
@@ 1 Introduction
The cornerstone of quantum estimation is the quantum Cramér-Rao bound (QCRB) [4, 3, 53, 54] which places a fundamental restriction on the precision in estimating unknown parameters of a quantum state. For one-dimensional parameters the bound is attainable in the limit of many copies, by measuring a specific observable called the symmetric logarithmic derivative, which is the quantum analogue of the classical score function. However, when dealing with complex models involving correlated states of many-body systems, optimal measurements may be hard to compute and implement in practice. Therefore, it is particularly important to devise realistic measurement schemes which allow the estimation of unknown parameters with close to optimal precision, by means of computationally efficient estimators.
@@
In this paper we provide a general measurement and data processing protocol for estimating an arbitrary dynamical parameter of a quantum Markov chain (QMC), which satisfies the above criteria. A QMC is a discrete time model of an open quantum system, in which the system interacts successively with a sequence of identically prepared “noise units” representing the environment, cf. Figure 1 a). The setup is similar to Haroche’s photon-box one-atom maser [55] and to that used in quantum collision models [56], and provides a physical mechanism for generating versatile many-body states such as matrix product states [57, 58] and finitely correlated states [59, 60]. By discretising time, QMCs can be used to model continuous-time dynamics of a Markovian open system coupled with Bosonic input-output channels [61, 62, 63].
@@
Figure 1: Basic elements of the pattern counting estimator. Panel a) A quantum Markov chain as a system interacting sequentially with the environment via a parameter dependent unitary $U_{\theta}$. The first stage estimator $\tilde{\theta}_{n}$ is obtained by performing a standard sequential measurement on the output and equating empirical and expected counts. Panel b) Post-processing the output using a coherent absorber. When system and absorber parameters match, the output is identical to the input (vacuum) Panel c) After the first estimation stage the absorber is fixed at a value $\theta_{\rm abs}=\tilde{\theta}_{n}-\delta_{n}$ where $\tilde{\theta}_{n}$ is the preliminary estimator and $\delta_{n}$ is the parameter shift required by the displaced-null measurement theory [1]. The output generated by the system and absorber dynamics with unitary $V_{\theta_{\rm abs}}U_{\theta}$ is measured sequentially in the standard basis. Panel d) Given a measurement trajectory, excitation patterns are identified as binary sequences starting and ending with a 11 separated by long sequences of 0s. The final estimator is a correction to the preliminary estimator which depends only on the total number of patterns $\sum_{\alpha}N_{\alpha,n}$, the QFI $f$ at $\tilde{\theta}_{n}$ and the displacement parameter $\tau_{n}$.
@@
For clarity, it is useful to distinguish between two mainstream approaches to parameter estimation in quantum open systems. In the setting of [64, 65, 18, 66, 67, 68, 69], the quantum system undergoes a noisy evolution depending on an unknown parameter, and the experimenter tries to extract information about the parameter by repeatedly applying instantaneous direct measurements and control operations while the system is evolving. In contrast, in this paper we adopt the setting commonly used in quantum optics and input-output theory [70, 71, 72, 73] where the experimenter does not have direct access to the system but can measure the output field of an environment channel coupled to the system. This allows the experimenter to track the conditional state of the system by means of stochastic filtering equations [74, 75, 76, 77, 78] and control it using feedback. As these techniques require full knowledge of the system’s dynamical parameters, it is important to devise tools for estimating such parameters from the stochastic trajectory of the measurement record.
Since the early works [79, 80], many aspects of continuous-time estimation have been investigated, including adaptive estimation [81, 82] filtering methods [83, 84, 85], Heisenberg scaling [86, 45, 87, 88], sensing with error correction [89], Bayesian estimation [90, 91, 92, 93, 94], quantum smoothing [95, 96, 97, 98], estimation of linear systems [99, 100, 101, 102], central limit and large deviations theory for trajectories [103, 104, 105, 106], concentration bounds for time averaged observables [107, 108], estimation with feedback control [109]. However, while the the ultimate precision limit can be expressed in terms of the output quantum Fisher information (QFI) [110, 111, 112, 113, 114, 100], and can in principle be attained by measuring a certain output observable, standard measurement protocols such as counting and homodyne typically do not achieve this limit. Therefore, attaining this bound with realistic measurements has been an important open problem in the field.
@@
Two recent papers [115, 2] have addressed this problem by introducing the idea of quantum post-processing of the output state using a quantum coherent absorber [116]. For a given QMC reference dynamics, the absorber takes the system’s output as its own input, and is characterised by the property that it “reverts” the action of the system so that the absorber’s output is a trivial product state (vacuum in continuous-time dynamics), cf. Figure 1 b). In the statistical estimation framework, the absorber is set to a particular reference parameter of the QMC dynamics, which is kept fixed throughout the protocol. A small deviation of the true parameter from this value will lead to non-trivial output statistics which can be used to estimate the deviation as illustrated in Figure 1 c). In [115], two of the present authors proved that the QCRB can be achieved by performing sequential, adaptive measurements on the output units, after the interaction with the coherent absorber. In addition, the adaptive measurement can be implemented efficiently in a Markovian fashion. The work [2] deals with the same problem but in the continuous-time setting, and proposes to perform a standard counting measurement (instead of an adaptive one) in conjunction with post-processing using the coherent absorber. In both papers the final estimator was computed from the measurement trajectory using the maximum likelihood method, which becomes computationally expensive for long trajectories.
@@
The strategy proposed here is similar to that of [2], albeit in discrete rather than continuous time, but strengthens it in several important aspects. Firstly, we employ the technique of displaced-null measurements [1] to provide a precise recipe for choosing the coherent absorber parameter. As we explain below, this is an important technical detail, as the intuitive choice of absorber parameter fails to achieve the QCRB in the limit of long trajectories. Secondly, our strategy employs a two step adaptive procedure which allows us to compute the final estimator as a simple linear transformation of the total number of “pattern counts” which can be easily extracted from the measurement trajectory, cf. Figure 1 d). This circumvents the computational cost associated with other estimators such as maximum likelihood, and the lack of theoretical results on their performance, specific to this context. Thirdly, we provide strong theoretical evidence that the final estimator achieves the QCRB in the limit of large times. This is based on a novel asymptotic representation of the output in terms of translationally invariant modes which obey canonical commutation relations and whose state is shown to satisfy the quantum local asymptotic normality property [117, 118, 119, 120]. This allows us to reduce the QMC estimation problem to a more familiar Gaussian estimation one, and cast the sequential output measurement as a counting measurement of the Gaussian modes.
@@
We now give a brief summary of the two stages estimation strategy proposed here and the related mathematical results. We consider a QMC whose dynamics depends on a one-dimensional parameter $\theta$ which we aim to estimate by measuring the output state produced after $n$ time steps. In the first stage we run the QMC dynamics with the unknown parameter $\theta$ for $1\ll\tilde{n}\ll n$ time steps and measure the noise units in a fixed basis, cf. Figure 1 a). From the total counts statistics we construct a rough estimator $\tilde{\theta}_{n}$ by matching the empirical frequency to its expected value. In general this estimator is neither optimal nor unbiased, but its mean square error has the standard $1/\tilde{n}$ scaling [108]. Its role is to reduce the range of the unknown parameter to a shrinking region, and allow the second stage measurement to be optimal in this region, see [1] for details in the case of independent samples. In the second stage we run the system and absorber QMC for the remaining $n-\tilde{n}$ time steps and measure the output in the standard basis. If the absorber parameter matched the true system parameter $\theta$, this measurement would produce a string of 0s (corresponding to no counts in continuous-time) cf. Figure 1 b). Therefore, it would seem natural to choose the absorber parameter to be $\tilde{\theta}_{n}$, our best guess at the unknown parameter $\theta$. However, this choice is unsuitable since for small deviations $\Delta_{n}=\theta-\tilde{\theta}_{n}$, the counting statistics depends quadratically on $\Delta_{n}$, which prevents the estimation of $\theta$ at standard $1/n$ rate. This non-identifiability issue is explained in detail in [1], which also provides the solution to this problem. We deliberately set the absorber parameter at $\theta_{\rm abs}=\tilde{\theta}_{n}-\delta_{n}$, which is away from the best guess by a small “displacement” $\delta_{n}\downarrow 0$ chosen to be larger than the uncertainty $|\Delta_{n}|$. This allows us to unambiguously identify $\theta$ from counts statistics. Stage two of the estimation procedure is illustrated in Figure 1 c).
We turn now to the question of estimating $\theta$ from the counts trajectory $\omega=(\omega_{1},\omega_{2},\dots,\omega_{n})$ of the second stage measurement (setting $n-\tilde{n}$ to $n$ for simplicity). Since $\theta-\theta_{\rm abs}$ is vanishingly small (even with the extra displacement), $\omega$ will typically contain a small number of $1$s while most of the outcomes will be 0, cf. Figure 1 d). This allows us split the trajectory into long sequences of 0s and in between them, binary “excitation patterns” starting and ending with a $1$. For each pattern
@@
$\alpha$ (e.g. $1$, $11$, $101$ etc.) we count the number of occurrences $N_{\alpha,n}$. In Theorem 3 we show that in the limit of large $n$ the counts $N_{\alpha,n}$ become independent Poisson variables whose intensities are $\lambda_{\alpha}u^{2}$ where $u=\sqrt{n}(\theta-\theta_{\rm abs})$ is the “local parameter” and $\lambda_{\alpha}$ is a model dependent coefficient which can be computed explicitly. Moreover the total Fisher information of the Poisson variables is equal to the output QFI, which shows that the pattern counts statistics capture the full information of the output state. Using this asymptotic behaviour, we construct a simple estimator $\hat{\theta}_{n}$ (cf. equations (25) and (24)) which is linear in the total pattern count, and we argue why it should achieve the QCRB in the limit of large $n$. To summarise, the two step procedure provides a computationally and statistically efficient estimation method which involves only standard basis measurements and a minimal amount of “quantum post-processing” implemented by the coherent absorber.
@@
For a more in-depth understanding of why the excitation pattern counts have asymptotically Poisson distributions, we refer to sections 4 and 5 where we develop a theory of translationally invariant modes (TIMs) of the output. These modes turn out to capture all statistical information about the unknown parameter, and can be measured simultaneously and optimally by performing the sequential standard output measurement. For each excitation pattern $\alpha=(\alpha_{1},\dots,\alpha_{k})\in\{0,1\}^{k}$ we define the creation operator $A^{*}_{\alpha}(n)$ on the output chain of length $n$. This consists of a running average
@@
where $\sigma^{\alpha}_{i}$ is the tensor product of the type $\sigma^{\alpha}=\sigma^{\alpha_{1}}\otimes\dots\otimes\sigma^{\alpha_{k}}$ where $\sigma^{0}=\mathbf{1}$ and $\sigma^{1}=\sigma^{+}=|1\rangle\langle 0|$, with first tensor acting on position $i$ of the output chain. In Proposition 1 and Corollary 1 we show that asymptotically with $n$, by applying the creation operators to the reference (vacuum) state $|0\rangle^{\otimes n}$, we obtain Fock-type states with different excitation pattern numbers. The creation and annihilation operators satisfy the Bosonic commutation relations with each excitation pattern being an independent mode. For large $n$, a Fock state is a superposition of basis states consisting of long sequences of 0s interspersed with the corresponding patterns appearing in any possible order. One of our key results, Theorem 2 shows that when the gap between system and absorber parameters scales as $\theta-\theta_{\rm abs}=u/\sqrt{n}$, the quadratures of the excitation pattern modes satisfy the Central Limit Theorem and the corresponding joint state is a product of coherent states whose amplitudes are linear in the local parameter $u$. The total QFI of this multimode coherent state is equal to the output QFI, showing that the TIMs contain all statistical information about the dynamics. We also prove separately, that the number operators of the TIMs have asymptotic Poisson distributions, as expected for a coherent state. Together with the result of Theorem 3, this completes a circle of ideas, which played a crucial role in formulating our estimation strategy. In a nutshell, when looking at the output from the perspective of the TIMs, one deals with a simple Gaussian estimation problem. Using the displaced-null method, we can achieve the QCRB by measuring the number operators of the TIMs and such a measurement can be implemented by simple sequential counting measurements followed by the extraction of pattern counts from the measurement trajectory $\omega$.
@@
The paper is organised as follows. In section 2 we give a brief review of quantum estimation theory and the displaced-null measurement technique developed in [1]. In section 3 we introduce the notion of QMC and the estimation problem, together with the idea of quantum post-processing using a coherent absorber. In section 4 we define the translationally invariant modes of the output and establish their Fock space properties. In section 5 we show that the restriction of the output state to the TIMs is a coherent state whose amplitude is linear in the local parameter and whose QFI is equal to the output QFI (cf. Theorem 2 and Corollary 2). In section 6 we establish that the excitation pattern counts obtained from the sequential output measurement have asymptotically Poisson distribution (cf. Theorem 3). In section 7 we formulate our measurement and estimation strategy and define the ”pattern counts” estimator. Finally in section 8 we present results of a simulation study confirming the earlier theoretical results.
2 Quantum estimation and the displaced null measurement technique